OpenStudy (amit69):

Say whether the l'Hospital's rule applies. If it does, use it to evaluate the given limit. If not, use some other method limit x tends -1 (x^2+5x+4)/(x^2+x)

OpenStudy (amit69):

I found out that L' Hospital rule applies but cannot evaluate the limit

OpenStudy (eliesaab):

This limit can be evaluated using two different methods \[ \frac{x^2+5 x+4}{x^2+x}=\frac{(x+1) (x+4)}{x (x+1)}=\frac{x+4}{x}\to -3 \] as x goes to -1

OpenStudy (eliesaab):

The other method is to use L'Hospital Rule \[ \lim_{x->-1} \frac{x^2+5 x+4}{x^2+x}=\lim_{x->-1}\frac {2x+5}{2x+1}=\frac 3{-1}=-3 \]

OpenStudy (3mar):

I thank you in person, @eliesaab! Really it is helpful for me! Thank you very much.

OpenStudy (eliesaab):

You are very welcome @3mar

OpenStudy (3mar):

Thank you for good explanation. You are Awesome, @eliesaab!